# GRAG: Graph Retrieval-Augmented Generation

### v2

@@ Abstract
Naive Retrieval-Augmented Generation (RAG) focuses on individual documents during retrieval and, as a result, falls short in handling networked documents which are very popular in many applications such as citation graphs, social media, and knowledge graphs. To overcome this limitation, we introduce Graph Retrieval-Augmented Generation (GRAG), which tackles the fundamental challenges in retrieving textual subgraphs and integrating the joint textual and topological information into Large Language Models (LLMs) to enhance its generation. To enable efficient textual subgraph retrieval, we propose a novel divide-and-conquer strategy that retrieves the optimal subgraph structure in linear time. To achieve graph context-aware generation, incorporate textual graphs into LLMs through two complementary views—the text view and the graph view—enabling LLMs to more effectively comprehend and utilize the graph context. Extensive experiments on graph reasoning benchmarks demonstrate that in scenarios requiring multi-hop reasoning on textual graphs, our GRAG approach significantly outperforms current state-of-the-art RAG methods.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in a variety of reasoning tasks, including on graph-based data (Hu et al., 2023b; Chen et al., 2024; Fatemi et al., 2023). However, LLMs themselves struggle with factual errors due to limitations in their training data and a lack of real-time knowledge (Mallen et al., 2023; Min et al., 2023). Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020; Ram et al., 2023; Gao et al., 2023), which integrates external data retrieval into the generative process, has been widely used to help LLMs access relevant information from external sources to generate more relevant answers and hence reduce factual errors (Tang and Yang, 2024). Naive RAG approaches focus solely on individual documents and retrieve relevant ones based on text similarity. However, real-world documents, such as social media postings, research papers, knowledge items, and product reviews, are typically not isolated but networked as textual graphs (He et al., 2023; Jin et al., 2023; Li et al., 2023). Importantly, such network information is typically crucial in both retrieving relevant documents and prompting LLMs for text generation (Yang et al., 2024; Tang and Yang, 2024). For example, research papers form a citation graph, so when a solar physicist wants to learn state-of-the-art techniques in solar flare prediction, paper mutual referencing links need to be considered to pursue comprehensive retrieval coverage and insightful technical evolution understanding of this research community (as shown in Figure 1). Similarly, social interaction among social media postings, entity relations in knowledge graphs, and purchasing relations in product review systems are indispensable when LLMs want to leverage these external data. So the question is how LLMs could harness this type of networked documents when performing RAG?
@@ 4.2 Textual Graph Augmented Generation
$=\prod_{i=1}^{r}p_{\theta,\phi_{1},\phi_{2},\phi_{3},\Phi}(y_{i}|y_{<i},[\mathbf{h}_{\hat{g}};\mathbf{h}_{T}]),$ (11)

### v3

@@ Abstract
Naive Retrieval-Augmented Generation (RAG) focuses on individual documents during retrieval and, as a result, falls short in handling networked documents which are very popular in many applications such as citation graphs, social media, and knowledge graphs. To overcome this limitation, we introduce Graph Retrieval-Augmented Generation (GRAG), which tackles the fundamental challenges in retrieving textual subgraphs and integrating the joint textual and topological information into Large Language Models (LLMs) to enhance its generation. To enable efficient textual subgraph retrieval, we propose a novel divide-and-conquer strategy that retrieves the optimal subgraph structure in linear time. To achieve graph context-aware generation, incorporate textual graphs into LLMs through two complementary views—the text view and the graph view—enabling LLMs to more effectively comprehend and utilize the graph context. Extensive experiments on graph reasoning benchmarks demonstrate that in scenarios requiring multi-hop reasoning on textual graphs, our GRAG approach significantly outperforms current state-of-the-art RAG methods. Our datasets as well as codes of GRAG are available at https://github.com/HuieL/GRAG.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in a variety of reasoning tasks, including on graph-based data (Hu et al., 2023b; Chen et al., 2024; Fatemi et al., 2023; Pan et al., 2024). However, LLMs themselves struggle with factual errors due to limitations in their training data and a lack of real-time knowledge (Mallen et al., 2023; Min et al., 2023). Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020; Ram et al., 2023; Gao et al., 2023; Yu et al., 2025), which integrates external data retrieval into the generative process, has been widely used to help LLMs access relevant information from external sources to generate more relevant answers and hence reduce factual errors (Tang and Yang, 2024). Naive RAG approaches focus solely on individual documents and retrieve relevant ones based on text similarity. However, real-world documents, such as social media postings, research papers, knowledge items, and product reviews, are typically not isolated but networked as textual graphs (He et al., 2023; Jin et al., 2023; Li et al., 2023; Hu et al., 2025). Importantly, such network information is typically crucial in both retrieving relevant documents and prompting LLMs for text generation (Yang et al., 2024; Tang and Yang, 2024). For example, research papers form a citation graph, so when a solar physicist wants to learn state-of-the-art techniques in solar flare prediction, paper mutual referencing links need to be considered to pursue comprehensive retrieval coverage and insightful technical evolution understanding of this research community (as shown in Figure 1). Similarly, social interaction among social media postings, entity relations in knowledge graphs, and purchasing relations in product review systems are indispensable when LLMs want to leverage these external data. So the question is how LLMs could harness this type of networked documents when performing RAG?
@@ 4.2 Textual Graph Augmented Generation
$=\prod_{i=1}^{n}p_{\theta,\phi_{1},\phi_{2},\phi_{3},\Phi}(y_{i}|y_{<i},[\mathbf{h}_{\hat{g}};\mathbf{h}_{T}]),$ (11)
@@ 6 Conclusion
7 Limitations
While GRAG provides an effective framework for graph context-aware generation, the efficiency of GRAG’s textual subgraph retrieval depends on the quality of the initial node ranking and pruning mechanism. In cases where graph structure or node importance is difficult to estimate, this could lead to suboptimal retrieval performance.
Acknowledgments
This work was supported by the National Science Foundation (NSF) Grant No. 2414115, No. 2403312, No. 2007716, No. 2007976, No. 1942594, No. 1907805, NIH R01AG089806, and NIH R01CA297856.